OpenStudy (anonymous):

Can someone help me make percentages in a chart? Please, I really need help!

OpenStudy (anonymous):

Suppose you want to show a percentage breakdown of average annual earthquakes in Alaska since 1900. Use the data below to create a pie graph. You will need to convert the number of earthquakes to a percentage of the total number in each case. Use the third column to record the percentages. Add a title to your pie graph. Upload your pie graph.

OpenStudy (anonymous):

I will make the pie graph, but I am having trouble finding the percentages...

OpenStudy (anonymous):

add them..90 + 20 + 5 + 1 = 116 earthquakes 4 to 5 ---- 90 ----90/116 = 0.775 = 77.5% or if you need to round 78% 5 to 6 ----20 -----20/116 = 0.172 = 17.2 or if rounded = 17% 6 to 7 ---5-------5/116 = .043 = 4.3% 7 or higher --1----1/116 = .0086 = 0.86%

OpenStudy (anonymous):

See, that's what I was thinking of, but I wasn't sure about doing it or not, like dividing everything. But okay, So now I understand and I can make the pie graph!

OpenStudy (anonymous):

or, like I said, if you need to round them.. 77.5 = 78% 17.2 = 17% 4.3 = 4% 0.86 = 1% -------total 100%

undeadknight26 (undeadknight26):

U got it Ricardo?

undeadknight26 (undeadknight26):

Kellie got it spot on though...so no point in asking that question...

OpenStudy (anonymous):

Yeah dude, I got it all!

OpenStudy (anonymous):

Here is the chart @kelliegirl33 I had to use an online chart maker that was suggested in the lesson. I think it came out accurate!